# Echo State Networks for Spatio-Temporal Area-Level Data

### v1

@@ 4.3 Forecasting Results
To assess model performance, we evaluate root mean square error (RMSE), continuous ranked probability score (CRPS), and interval score (IS). Notably, CRPS evaluates the forecasting performance over the entire distribution, making it a useful measure of uncertainty quantification. In this study, we employ the empirical form of CRPS as outlined by Zamo and Naveau, (2018).
As shown in Table 1, the AESN outperforms all other models in both short-lead (3-step ahead) and long-lead (12-step ahead) forecasts, achieving the lowest RMSE, CRPS, and IS, indicating superior accuracy and uncertainty quantification. ESN with EOF also demonstrates strong performance, particularly with lower errors compared to the standard ESN. In contrast, the DLM struggles significantly in long-lead forecasts, with an extremely high RMSE and IS, making it the least effective model for long-lead forecasting.
@@
RMSE CRPS IS RMSE CRPS IS
DLM 1927.990 243.044 4172.548 27418371.293 251.873 1275719.846
ESN 477.667 166.814 1359.580 675.250 205.312 2391.955
ESN with EOF 326.096 103.508 1135.795 477.667 166.814 1359.580
AESN 202.307 60.429 561.487 332.766 91.588 908.191
@@ 5 Discussion
In this paper, we introduce a novel approach that combines random representation and GCNs to adeptly handle the spatial structure in ESNs. The efficiency of our method stems from the nature of the random representation, where most parameters are randomly sampled and then fixed, reducing the computational overhead while improving model performance. We demonstrate that our model significantly outperforms traditional DLMs and standard ESN approaches, including those enhanced with EOFs using the Eurostat tourism occupancy dataset. Our results show that the model not only captures the long-lead temporal dynamics effectively but also accurately reflects the intricate spatial dynamics within the data. This capability makes it exceptionally useful for understanding and predicting complex patterns in regional tourism activities, thereby aiding stakeholders in making more precise and effective policies and interventions.

### v2

@@ 4.3 Forecasting Results
To assess model performance, we evaluate root mean square error (RMSE), continuous ranked probability score (CRPS), and interval score (IS). Notably, CRPS evaluates the forecasting performance over the entire distribution, making it a useful measure of uncertainty quantification. In this study, we employ the empirical form of CRPS as outlined by zamo2018estimation.
As shown in Table 2, the AESN outperforms all other models in both short-lead (3-step ahead) and long-lead (12-step ahead) forecasts, achieving the lowest RMSE, CRPS, and IS, indicating superior accuracy and uncertainty quantification. ESN with EOF also demonstrates strong performance, particularly with lower errors compared to the standard ESN. In contrast, the DLM struggles significantly in long-lead forecasts, with an extremely high RMSE and IS, making it the least effective model for long-lead forecasting. In terms of runtime, all ESN-based models are clearly more efficient than the DLM model. Among them, the ESN with EOF is the fastest model, which is expected since it only considers the first eight principal components, as discussed in Section 4.2. We note that AESN achieves significant improvements in modeling area-level data with only minimal computational cost. In addition, the 3-step forecasts are slower because they require model updates whenever new data becomes available after each 3-step interval. In practice, runtime can also be affected by the choice of hyperparameters. For example, when the analysis requires a small value of $K$, AESN can be as fast as ESN.
@@
RMSE CRPS IS Time RMSE CRPS IS Time
DLM 1927.990 243.044 4172.548 1.57 hours 27418371.293 251.873 1275719.846 24.91 mins
ESN 477.667 166.814 1359.580 4.68 mins 675.250 205.312 2391.955 1.23 mins
ESN with EOF 326.096 103.508 1135.795 3.90 mins 477.667 166.814 1359.580 0.93 mins
AESN 202.307 60.429 561.487 5.52 mins 332.766 91.588 908.191 2.21 mins
@@ 5 Discussion
In this paper, we introduce a novel approach that combines random representation and GCNs to adeptly handle the spatial structure in ESNs. The efficiency of our method stems from the nature of the random representation, where most parameters are randomly sampled and then fixed, reducing the computational overhead while improving model performance. We demonstrate that our model significantly outperforms traditional DLMs and standard ESN approaches, including those enhanced with EOFs using the Eurostat tourism occupancy dataset. Our results show that the model not only captures the long-lead temporal dynamics effectively but also accurately reflects the intricate spatial dynamics within the data. This capability makes it exceptionally useful for understanding and predicting complex patterns in regional tourism activities, thereby aiding stakeholders in making more precise and effective policies and interventions. Finally, there are several potential directions for future research. First, the current implementation of AESN considers the full graph built on area-level data. Future research should explore training AESN using a mini-batch approach by constructing meaningful subgraphs that account for correlations between subgraphs. Second, missing data is quite common in official statistics due to disclosure limitations and the high costs of conducting surveys at fine geographical levels. Currently, neither the ensemble ESN nor AESN has a mechanism to account for missingness. Incorporating imputation methods could be a valuable direction for future research.